# Time-Series Foundation Model for Value-at-Risk

### v1

@@
11footnotetext: Tampere University, Department of Computing Sciences, Financial Computing and Data Analytics Group, Kalevantie 4, Tampere, 33100, Finland. Email: anubha.goel@tuni.fi, juho.kanniainen@tuni.fi22footnotetext: Department of Mathematics, Indian Institute of Technology Ropar, Rupnagar, Punjab, 140001, India. Email: puneet.pasricha@iitrpr.ac.in33footnotetext: Corresponding author: Juho Kanniainen
Time-Series Foundation Model for Value-at-Risk
Anubha Goel*    Puneet Pasricha    Juho Kanniainen*
(October 15, 2024)
This study is the first to explore the application of a time-series foundation model for VaR estimation. Foundation models, pre-trained on vast and varied datasets, can be used in a zero-shot setting with relatively minimal data or further improved through finetuning. We compare the performance of Google’s model, called TimesFM, against conventional parametric and non-parametric models, including GARCH, Generalized Autoregressive Score (GAS), and empirical quantile estimates, using daily returns from the S&P 100 index and its constituents over 19 years. Our backtesting results indicate that, in terms of the actual-over-expected ratio, the fine-tuned TimesFM model consistently outperforms traditional methods. Regarding the quantile score loss function, it achieves performance comparable to the best econometric approach, the GAS model. Overall, the foundation model is either the best or among the top performers in forecasting VaR across the 0.01, 0.025, 0.05, and 0.1 VaR levels. We also found that fine-tuning significantly improves the results, and the model should not be used in zero-shot settings. Overall, foundation models can provide completely alternative approaches to traditional econometric methods, yet there are challenges to be tackled.
Keywords: Value at Risk, Foundation Models, TimesFm, Time Series Forecasting
JEL classification: C5, G17
MSC: 91G70, 68T05
@@ 1 Introduction
While the application of Machine Learning (ML) in various fields in financial market research has gained significant attention—such as in asset pricing, volatility forecasting, optimal hedging, and limit order book dynamics (Gu et al., , 2020; Christensen et al., , 2023; Tran et al., , 2018; Mikkilä and Kanniainen, , 2023)—the risk management literature still largely relies on traditional econometric techniques, especially in the area of Value-at-Risk (VaR). One reason for this may be that Deep Learning (DL) models require substantial amounts of data, which poses a significant limitation when applying them to calculate risk metrics for individual stocks. VaR estimations are often based on daily return observations (for example, see González-Rivera et al., , 2004; Patton et al., , 2019; Taylor, , 2019, based solely on historical daily returns), and even 20 years of data would yield only around 5,000 observations. This is quite limited, especially considering that DL models often contain millions of parameters that need to be learned.
@@
However, a solution to this problem is currently underway. Recently, machine learning has experienced a paradigm shift with the emergence of foundation models, which are large-scale, general-purpose pre-trained and often fine-tunable models on diverse datasets spanning various data distributions. Although foundation models are built on standard deep learning and transfer learning principles, their scale leads to the emergence of new capabilities (Bommasani et al., , 2021).
In 2023–2024, following major advancements in large language models (LLMs), several pre-trained time-series foundation models have emerged, including Google’s TimesFM (Das et al., , 2024), Time-GPT (Garza and Mergenthaler-Canseco, , 2024), LagLlama (Rasul et al., , 2023). These time-series forecasting models are pre-trained on vast datasets of billions of real-world time points and can be used with or without fine-tuning, requiring only relatively small datasets, with just a few hundred observations. Compared to conventional dedicated supervised ML models, large- foundational AI models are pre-trained using large and diverse datasets. enable new applications more quickly and cost-effectively, and support a wider range of AI applications. The word ‘foundational’ emphasizes their flexibility, as they can be adapted or customized for different tasks or domains with minimal additional training.
@@
In this paper, we implement a pre-trained time-series foundation model, specifically Google’s TimesFM***The authors of Google’s TimesFM empirically demonstrated that their model outperformed TimeGPT in forecasting accuracy, which is why we chose to use TimesFM., for Value-at-Risk (VaR) estimation and compare it with existing state-of-the-art parametric and semi-parametric approaches, including GARCH model, Generalized Autoregressive Score (GAS) one-factor model, and empirical quantile estimates. The key question is whether the pre-trained foundation model can outperform these existing econometric methods. This paper could employ alternative foundation models as several have been published recently, and more are on the way. However, we decided to focus on the use TimesFM for two reasons: First, it is an open-source and free-of-chargehttps://github.com/google-research/timesfm., making our results completely replicableThe codes of this paper can be found at https://github.com/Anubha0812/TimesFM-for-Value-at-Risk..
Second, our goal is not to find the best possible foundation model among those available, as models are likely to be modified and entirely new ones published on a monthly basis. Rather, we aim to conduct an extensive empirical analysis to determine whether classical parametric and semi-parametric approaches can be outperformed by a purely data-driven foundation model. This is a more fundamental question than focusing on which foundation model is superior right now, as that will inevitably continue to evolve.
@@
In this paper, we conduct an extensive empirical evaluation of the performance of the foundation time-series model with daily return data on the 91 constituents of the S&P 100 index over 19 years. For each constituent, we calculate 1-day VaR on a daily basis. The foundation model is applied both with and without fine-tuning to understand how much the results improve with fine-tuning, which has certain computational cost. The fine-tuning is performed on a daily basis using a rolling window technique.
Our main finding is that the fine-tuned time-series foundation model outperforms existing econometric approaches for VaR estimation. This result is robust across various backtesting methods, and its reliability is further strengthened by the fact that its performance is, on average, superior across the 91 constituents of the S&P 100. We consider this as a remarkable result on the performance of Google’s TimesFM model, especially this model is not explicitely designed for this purpose. Moreover, the TimeFM was not pre-trained using the data used in this paper.§§§TimesFM was trained using Google trends, Wiki Pageviews, Synthetic Data, and Electricty, Traffic, and Weather datasets used in M4 competition (Makridakis et al., , 2020; Spiliotis et al., , 2020). Further, we demonstrate that two years of daily return data is sufficient to achieve accurate results through fine-tuning, indicating that the amount of data required for successful fine-tuning is not an issue. Secondly, while fine-tuning clearly offers advantages, we are somewhat surprised to find that the foundation model performs quite well even without fine-tuning. These findings have several implications beyond Value-at-Risk estimation techniques, which, along with the limitations of our analysis, will be discussed at the end of the paper in Section 6.
@@
The paper is structured as follows: In Section 2, we describe the time-series foundation model, Google’s TimesFM, and the benchmark models against which it is compared. Section 3 details the data and experimental setup for the models and fine-tuning specifications for VaR estimation. The evaluation metrics and backtesting procedures are outlined in Section 4. Section 5 presents a thorough analysis of the empirical results, examining the performance of the foundation model relative to traditional econometric approaches. Finally, in Section 6, we discuss the broader implications of our findings, exploring the potential opportunities and challenges that foundation models present for econometrics, particularly in the context of financial risk management.
@@ 2.1 Time-Series Foundation Models
In the case of time-series forecasting, these models share similarities with large language models (LLMs), as they rely on transformer-based architectures. While LLMs are trained to predict the next or missing part of a text sequence, foundational time-series models process continuous time points as individual tokens. This approach allows them to handle time-series data step by step. Similarly to Large Language Models, time-series foundation models utilize transformers to capture long-range dependencies and patterns within sequential data.
The TimesFM model (Das et al., , 2024) is a decoder-only transformer architecture, similar to GPT models. In contrast to traditional encoder-decoder models, where the encoder processes the input sequence and the decoder generates output based on the encoded input, decoder-only models like TimesFM the decoder-only model focuses solely on output generation. processing the input through successive decoder layers. This allows for faster training and inference times compared to encoder-decoder models (see, for example Krishnan Suresh et al., , 2024). Decoder-only models like are highly advantageous for zero-shot forecasting, where minimal fine-tuning is required. For that reason, in this paper, we use TimesFM as a foundation model as our aim is to test the performance of foundation models with small-sample fine-tuning properties. Moreover, in contrast to TimeGPT, it is open-source solution with proper documentation (Das et al., , 2024). On the other hand, the advantage of the encoder-decoder architecture, which is used, for example, in TimeGPT (Garza and Mergenthaler-Canseco, , 2024), is that it can handle more complex inputs and exogenous variables. However, in this paper, we use univariate and well-structured time-series data on returns, and VaR is calculated separately for each security.
Compared to the state-of-the-art LLMs, TimesFM time-series foundation model is much smaller in both parameter size with 200M parameters and pretraining data size with around 100B timepoints. However, these settings allow it to achieve accurate zero-shot forecasts, meaning it generate forecasts from unseen datasets.
@@ 2.2 Benchmark Models for VaR
    One of the most widely used approaches for VaR estimation is the rolling window method, which computes the risk measure based on historical data over a fixed time window. This method is defined as follows:
@@
  • ARMA-GARCH Models (Angelidis et al., , 2004):
    To better capture the time-varying properties of financial returns, we employ ARMA-GARCH models. These models estimate the conditional mean and variance of returns, incorporating various distributional assumptions for the standardized residuals. The model is specified as:
    $r_{t}=\mu_{t}+\sigma_{t}\eta_{t},$
    $\eta_{t}\sim\text{iid }\mathbb{G}_{\eta}(0,1),$
    In this framework, $\mu_{t}$ represents the conditional mean, typically modeled using an ARMA process, while $\sigma_{t}^{2}$ denotes the conditional variance, modeled using a GARCH process. The standardized residuals $\eta_{t}$ follow a distribution $\mathbb{G}_{\eta}$ with mean zero and variance one. The VaR for these models is calculated as follows:
    $v_{t}=\mu_{t}+\sigma_{t}a_{\alpha},\quad\text{where }a_{\alpha}=\mathbb{G}_{\eta}^{-1}(\alpha)$
@@
      $\eta_{t}\sim\text{iid Skew }t(0,1,\nu,\lambda),$
@@
    3. (iii)
      Empirical Distribution Function: As a nonparametric alternative, we use the Empirical Distribution Function (EDF) method (Manganelli and Engle, , 2004). This approach estimates the distribution of standardized residuals $\eta_{t}$ from historical data, avoiding any parametric assumptions.
  • One factor Gas model (Patton et al., , 2019): This model belongs to the class of CAViaR models, where the evolution of the risk measure is directly parameterized without the need to specify a full conditional distribution for asset returns. The evolution of VaR for the Gas 1 factor model is given as follows:
@@
    where $a>0$ is a scaling parameter and $\kappa_{t}$ represents the dynamic risk factor driving the VaR process. The evolution of $\kappa_{t}$ is given by the updating mechanism:
    $\kappa_{t}=\omega+\beta\kappa_{t-1}+\gamma\frac{1}{\exp(\kappa_{t-1})}\left(\frac{1}{\alpha}\mathbf{1}\{r_{t-1}\leq a\exp(\kappa_{t-1})\}r_{t-1}-b\exp(\kappa_{t-1})\right).$
@@ 3 Data and Experiments
We now apply the models discussed in the last section to the forecasting of VaR for daily returns on the S&P 100 index along with its 91 constituents. We obtained the data from the Thomson Reuters Datastream, and our study period covers approximately 19 years from January 2005 to September 30, 2023, resulting in 4,876 observations. The data period was selected to maximize the availability of constituents, resulting in 91 stocks in our datasetThe S&P100 Global Index typically comprises 100 companies; however, as of October 1, 2023, it included 102 stocks (as Roche Holding AG and GOOGLE have two classes of shares). However, due to the unavailability of data for certain stocks, specifically Synlogic Inc (SYBX UR Equity), Alphabet Inc (GOOG UW Equity), Seven & i Holdings Co Ltd (3382 JT Equity), Texas Instruments Inc (TXN UW Equity), Shell PLC (SHEL LN Equity), Philip Morris International Inc (PM UN Equity), Mastercard Inc (MA UN Equity), Broadcom Inc (AVGO UW Equity), Engie SA (ENGI FP Equity), PepsiCo Inc (PEP UW Equity), and Honeywell International Inc (HON UW Equity)—the analysis was conducted on the remaining 91 constituents. The daily return for $i$-th stock at $t$-th day is calculated as $r_{it}=\frac{p_{it}-p_{it-1}}{p_{it-1}};$ $i=1,\ldots,n,\ t=1,\ldots,T,$ where $p_{it}$ and $p_{it-1}$ are respectively, the closing prices at $t$-th and $(t-1)$-th day. We divide the data into two consecutive non-overlapping samples while maintaining the temporal ordering of the data. We use the first ten years of data, i.e., January 2005 to December 2014, for estimating model parameters and the remaining 9 years, i.e., January 2015 to September 2023, for testing the performance of various models.

### v2

@@
Abstract
This study is the first to explore the application of a time-series foundation model for Value-at-Risk (VaR) estimation. Foundation models, pre-trained on vast and varied datasets, can be used in a zero-shot setting with relatively minimal data or further improved through finetuning. We compare the performance of Google’s model, called TimesFM, against conventional parametric and non-parametric models, including GARCH, Generalized Autoregressive Score (GAS), and empirical quantile estimates, using daily returns from the S&P 100 index and its constituents over 19 years. Our backtesting results indicate that in terms of the actual-over-expected ratio the fine-tuned TimesFM model consistently outperforms traditional methods. Regarding the quantile score loss function, it achieves performance comparable to the best econometric approach, the GAS model. Overall, the foundation model is either the best or among the top performers in forecasting VaR across the 0.01, 0.025, 0.05, and 0.1 VaR levels. We also found that fine-tuning significantly improves the results, and the model should not be used in zero-shot settings.
keywords:
Value at Risk , Foundation Models , TimesFM , Time series forecasting
journal: SSRN
[…]
[inst1]organization=Tampere University, Department of Computing Sciences, Financial Computing and Data Analytics Group,addressline=Kalevantie 4, city=Tampere, postcode=33100, country=Finland
[…]
[inst2]organization=IIT Ropar, Department of Mathematics,addressline=Ropar, city=Punjab, postcode=140001, country=India
@@ 1 Introduction
While the application of Machine Learning (ML) in various fields in financial market research has gained significant attention—such as in asset pricing, volatility forecasting, limit order book dynamics, and optimal hedging (Gu et al., 2020; Christensen et al., 2023; Tran et al., 2018; Mikkilä and Kanniainen, 2023)—the risk management literature still largely relies on traditional econometric techniques, especially in the area of Value-at-Risk (VaR). One reason for this may be that Deep Learning (DL) models require substantial amounts of data, which poses a significant limitation when applying them to calculate risk metrics for individual stocks. VaR estimations are often based on daily return observations (for example, see González-Rivera et al., 2004; Patton et al., 2019; Taylor, 2019, based solely on historical daily returns), and even 20 years of data would yield only around 5,000 observations. This is quite limited, especially considering that DL models often contain millions of parameters that need to be learned.
@@
However, a solution to this problem is currently underway. Recently, machine learning has experienced a paradigm shift with the emergence of foundation models, which are large-scale, general-purpose pre-trained and often fine-tunable models on diverse datasets spanning various data distributions. Although foundation models are built on standard deep learning and transfer learning principles, their scale leads to the emergence of new capabilities (Bommasani et al., 2021). In 2023–2024, following major advancements in large language models (LLMs), several pre-trained time-series foundation models have emerged, including Google’s TimesFM (Das et al., 2024), Time-GPT (Garza and Mergenthaler-Canseco, 2024), and LagLlama (Rasul et al., 2023).
AI-driven advancements in financial risk modeling have opened doors to new methodologies for assessing complex financial data (Tran et al., 2019; Guijarro-Ordonez et al., 2021; Chen et al., 2024). Historically, methods like GARCH-based models have been effective for capturing dependencies and volatility clustering in market data Hoga and Demetrescu (2023). Recent approaches, however, such as Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), and Conditional GANs (CGANs), extend these capabilities by generating realistic synthetic data for distribution forecasting (Nareklishvili et al., 2023; Athey et al., 2021; Hofert et al., 2022). These generative models allow for flexible, non-parametric modeling of data structures that were once challenging to approximate with traditional econometric tools. Ericson et al. (2024) provide a comprehensive comparative review, highlighting that deep generative methods outperform traditional benchmarks in capturing complex financial dynamics for VaR estimation. These advancements underscore the potential of AI-driven methods in risk forecasting. Our approach diverges by utilizing a time-series foundation model, pre-trained on extensive and various data sets, to predict VaR. Compared to traditional supervised ML models, which are trained for specific tasks, these time-series forecasting models are pre-trained on vast, diverse datasets containing billions of real-world time points. A single pre-trained model can be applied across a wide range of domains, from traffic forecasting to stock price predictions, offering quick and cost-effective solutions. The word ‘foundational’ emphasizes their flexibility, as they can be adapted for different tasks or domains with minimal additional training. The foundation models can be used with or without fine-tuning.
@@
We implement a pre-trained time-series foundation model, specifically Google’s TimesFM, for VaR estimation and compare it with existing state-of-the-art parametric and semi-parametric approaches, including GARCH model, Generalized Autoregressive Score (GAS) one-factor model, and empirical quantile estimates. The key question is whether the pre-trained foundation model yields results comparable to, or even capable of outperforming, existing econometric methods. Additionally, this paper seeks to determine how much fine-tuning enhances accuracy in the context of VaR predictions. This paper could employ alternative foundation models as several have been published recently, and more are on the way. However, we decided to focus on the use TimesFM because it is an open-source [^1] https://github.com/google-research/TimesFM., modifiable, and free-of-charge, [^2] As an additional note, the authors of Google’s TimesFM empirically demonstrated that their model outperformed TimeGPT in forecasting accuracy. making our results completely replicable. [^3] The codes of this paper can be found at https://github.com/Anubha0812/TimesFM-for-Value-at-Risk. Moreover, our goal is not to find the best possible foundation model among those available, as they are likely to be modified and entirely new ones can be published on a monthly basis, but rather to address more fundamental question if the state-of-the-art parametric and semi-parametric approaches can be outperformed.
@@
We conduct an extensive empirical evaluation of the performance of the foundation time-series model with daily return data on the 91 constituents of the S&P 100 index over 19 years. For each constituent, we calculate 1-day VaR on a daily basis. The foundation model is applied both with and without fine-tuning to understand how much the results improve with fine-tuning, which has certain computational cost. The fine-tuning is performed on a daily basis using a rolling window technique.
Our main finding is that, in terms of the actual-over-expected ratio, the fine-tuned time-series foundation model outperforms existing econometric approaches for VaR estimation. Specifically, the fine-tuned foundation model consistently demonstrates superior performance, on average, across all 91 constituents of the S&P 100 included in the analysis. Regarding the quantile score loss function, which was the second backtesting metric, the fine-tuned foundation model achieves performance comparable to the best econometric approach, the GAS model. Overall, the foundation model is either the top performer or among the best in forecasting VaR across the 0.01, 0.025, 0.05, and 0.1 VaR levels. Moreover, we found that fine-tuning is practically necessary for the foundation model; using it as pre-trained in zero-shot settings results in clearly worse performance compared to the best econometric approaches. We consider this as a remarkable result on the performance of Google’s TimesFM model as this model is not explicitly designed for this purpose and was not even pre-trained with stock return data. This problem is statistically challenging due to its focus on heteroscedastic processes with time-varying conditional volatility. In this light, the paper is valuable not only for financial risk management but also as a benchmark to test AI time-series models against state-of-the-art statistical approaches on a highly complex issue.
@@
The paper is structured as follows: In Section 2, we describe the time-series foundation model, Google’s TimesFM, and the benchmark models against which it is compared. Section 3 details the data and experimental setup for the models and fine-tuning specifications for VaR estimation. The evaluation metrics and backtesting procedures are outlined in Section 4. Section 5 presents an analysis of the empirical results, examining the performance of the foundation model relative to traditional econometric approaches. Finally, the findings have several implications beyond Value-at-Risk estimation techniques, which, along with the limitations of our analysis, will be discussed at the end of the paper in Section 6.
@@ 2.1 Time-Series Foundation Models
The time-series foundation models share similarities with large language models (LLMs) as they rely on transformer-based architectures to capture long-range dependencies and patterns within sequential data. While LLMs are trained to predict the next or missing part of a text sequence, foundational time-series models process continuous time points as individual tokens. This approach allows them to handle time-series data step by step.
TimesFM model (Das et al., 2024) is a decoder-only transformer architecture, similar to GPT models. In contrast to encoder-decoder models, where the encoder processes the input sequence and the decoder generates output based on the encoded input, decoder-only models focuse solely on output generation processing the input through successive decoder layers. This allows for faster training and inference times compared to encoder-decoder models. Decoder-only models like are highly advantageous for zero-shot forecasting, where minimal fine-tuning is required. On the other hand, the advantage of the encoder-decoder architecture, which is used, for example, in TimeGPT (Garza and Mergenthaler-Canseco, 2024), is that it can handle more complex inputs and exogenous variables. However, in this paper, we use univariate and well-structured time-series data on returns, and VaR is calculated separately for each security, and for that reason the encoder component does not add significant value.
The architecture of TimesFM is optimized for long-horizon forecasting through the use of a patching strategy, where the time-series data is broken into patches. These patches function similarly to tokens in language models, significantly improving efficiency. Unlike some autoregressive models that predict one step at a time, TimesFM predicts longer output patches at once, reducing the number of necessary auto-regressive steps. This makes the model particularly well-suited for handling varying context lengths, prediction lengths, and temporal granularities during inference.
In pretraining, TimesFM was trained on a vast corpus comprising both real-world time-series data such as Google trends, Wiki Pageviews, Synthetic Data, and Electricity, Traffic, and Weather datasets used in M4 competition (Makridakis et al., 2020; Spiliotis et al., 2020). However, financial data was not used, and this aspect makes the application of the present paper particularly interesting. This corpus, although smaller in scale compared to text-based models, spans around 100 billion time points, providing the diversity needed for the model to generalize well across domains and forecasting tasks. Compared to the state-of-the-art LLMs, TimesFM time-series foundation model is much smaller, with 200M parameters and pretraining data consisting of around 100B timepoints. Despite its smaller scale, it was reported to achieve relatively high zero-shot forecasting accuracy
@@ 2.2 Benchmark Models for VaR
    One of the most widely used approaches for VaR estimation is the rolling window method, which computes the risk measure based on historical data over a fixed time window:
@@
  • 2.
    ARMA-GARCH Models (Angelidis et al., 2004):
    To better capture the time-varying properties of financial returns, we employ ARMA-GARCH models. These models estimate the conditional mean and variance of returns, incorporating various distributional assumptions for the standardized residuals. The model is specified as
    $r_{t}=\mu_{t}+\sigma_{t}\eta_{t},\text{\quad}\eta_{t}\sim\text{iid }\mathbb{G}_{\eta}(0,1).$
    In this framework, $\mu_{t}$ represents the conditional mean, typically modeled using an ARMA process, while $\sigma_{t}^{2}$ denotes the conditional variance, modeled using a GARCH process. The standardized residuals $\eta_{t}$ follow a distribution $\mathbb{G}_{\eta}$ with mean zero and variance one. The VaR for these models is calculated as
    $v_{t}=\mu_{t}+\sigma_{t}a_{\alpha},\quad\text{where }a_{\alpha}=\mathbb{G}_{\eta}^{-1}(\alpha).$
@@
      $\eta_{t}\sim\text{iid Skew }t(0,1,\nu,\lambda).$
@@
    3. (c)
      Empirical Distribution Function: As a nonparametric alternative, we use the Empirical Distribution Function (EDF) method (Manganelli and Engle, 2004). This approach estimates the distribution of standardized residuals $\eta_{t}$ from historical data, avoiding any parametric assumptions.
  • 3.
    One factor Gas model (Patton et al., 2019): This model belongs to the class of CAViaR models (Taylor, 2005), where the evolution of the risk measure is directly parameterized without the need to specify a full conditional distribution for asset returns. [^4] GAS model was originally proposed in Creal et al. (2013). Patton et al. (2019) proposed a dynamic GAS model employing the class of loss functions proposed in Fissler and Ziegel (2016), called FZ loss functions. In this paper, we adopt the GAS model from Section 2.3 in Patton et al. (2019). The evolution of VaR for the Gas 1 factor model is given
@@
    where $a<0$ is a scaling parameter and $\kappa_{t}$ represents the dynamic risk factor driving the VaR process. The evolution of $\kappa_{t}$ is given by the updating mechanism
    $\begin{split}\kappa_{t}&=\omega+\beta\kappa_{t-1}\\ &+\gamma\frac{1}{b\exp(\kappa_{t-1})}\left(\frac{1}{\alpha}\mathbf{1}\{r_{t-1}\leq a\exp(\kappa_{t-1})\}r_{t-1}-b\exp(\kappa_{t-1})\right),\end{split}$
    where $b<a<0<\gamma$ and we set $\omega=0$ as in Patton et al. (2019)
@@ 3 Data and Experiments
We now apply the models discussed in the last section to the forecasting of VaR for daily returns on the S&P 100 index along with its 91 constituents. The data used in our analysis was obtained from Thomson Reuters Datastream, covering a period of approximately 19 years, from January 2005 to September 30, 2023, resulting in 4,876 observations. Throughout this period, data on every day was available for 91 stocks. [^5] The S&P100 Global Index typically comprises 100 companies; however, as of October 1, 2023, it included 102 stocks (as Roche Holding AG and GOOGLE have two classes of shares). However, due to the unavailability of data for certain stocks, specifically Synlogic Inc (SYBX UR Equity), Alphabet Inc (GOOG UW Equity), Seven & i Holdings Co Ltd (3382 JT Equity), Texas Instruments Inc (TXN UW Equity), Shell PLC (SHEL LN Equity), Philip Morris International Inc (PM UN Equity), Mastercard Inc (MA UN Equity), Broadcom Inc (AVGO UW Equity), Engie SA (ENGI FP Equity), PepsiCo Inc (PEP UW Equity), and Honeywell International Inc (HON UW Equity)—the analysis was conducted on the remaining 91 constituents. The daily return for $i$-th stock at $t$-th day is calculated as $r_{it}=\frac{p_{it}-p_{it-1}}{p_{it-1}};$ $i=1,\ldots,n,\ t=1,\ldots,T,$ where $p_{it}$ and $p_{it-1}$ are respectively, the closing prices at $t$-th and $(t-1)$-th day. We divide the data into two consecutive non-overlapping samples while maintaining the temporal ordering of the data. We use the first ten years of data, i.e., January 2005 to December 2014, for estimating model parameters and the remaining 9 years, i.e., January 2015 to September 2023, for testing the performance of various models.